Consider the utility function of an individual given by u(x) - ln(x - 10, 000). His total wealth is $270,000 of which S170,000 is the worth of his house. There is 10% probability that his house may be destroyed by fire. (a) What is the risk attitude of this person? (10%) (b) Calculate the insurance premium. fair premium and risk premium. (1596) (c) What is the relationship between the insurance premium, fair premium, and 2.
Answer #1

Answer:

Utility function U=ln(X-10000)

A)

Marginal utility of wealth=dU/dX=1/(X-10000)

Since as the wealth increases the marginal utility decreases because marginal utility and wealth has inverse relationship. So the person is risk averse.

B)

Total Wealth T= $2,70,000

Worth of House H=$ 1,70,000

If there is no Fire

Total wealth = X=$2,70,000

If there is Fire

Total Wealth = X=T-H=2,70,000-1,70,000=$1,00,000

So expected utility E(U)=(1-Probability of fire)*U(2,70,000)+ Probability of Fire*U(1,00,000)

E(U)=90%*ln(270000-10000)+10%*ln(100000-10000)=11.22+1.14=12.36

Wealth for Expected Utility = e^(12.36)+10000=$243281.23

Insurance premium=270000-243281.23=$26718.77

fair Premium =10%*1,70,000=$17000

So risk Premium =Insurance Premium - Fair Premium =26718.77-17000=$9718.77

C)

Risk Premium =Insurance Premium - Fair Premium